# Fixed-Point RNNs: From Diagonal to Dense in a Few Iterations

### v1

@@ Appendix D Low-Rank Expressiveness
Following the notation and methodology of (Cirone et al., 2024b)[B.4] ), this can be written in terms of the Signature as
@@
It follows directly from (15) that any linear readout of $Y_{t}$ can be represented as a series in signature terms. As a result, these systems are fundamentally restricted to learning functions that closely approximate these convergent series.
@@
In our setting $\frac{1}{N}\langle A_{I}Y_{0},A_{J}Y_{0}\rangle_{\mathbb{R}^{N}}$ corresponds to a product graph $G_{I,J}$ corresponding to a ladder having $2|I|+2|J|$ edges as shown in figure 8. We can then use (Cirone et al., 2024a)[Prop. 2] to compute the square of the $L^{2}$ norm in equation (17), the only difference from the dense case is that half of the vertices (excluding the ”middle” one) correspond to a space of dimension $r_{N}$ while the rest to the standard $N$.
@@
Since the total number of admissible pairings of $G_{I,J}\sqcup G_{I,J}$ is $\bigl{(}4(|I|+|J|)\bigr{)}!!$!, we conclude that equation (17) holds with $\kappa=4$ and $o(1):=\mathcal{O}\bigl{(}\frac{1}{\sqrt{N\wedge r_{N}}}\bigr{)}$.
@@
Remark D.4.
While the proof crucially uses the assumption $r_{N}\to\infty$ as $N\to\infty$, at the same time we have not provided an argument against $r_{N}$ not diverging. In figure 9 we present a counterexample, showing that if $r_{N}$ does not diverge then the asymptotics differ from the dense ones, in particular some symmetries are ”lost”, impossible to recover due to unavoidable noise.
Figure 9: Admissible pairing different from the ”identity” paring, but still leading to maximal asymptotic scaling in the bounded $r_{N}$ case. Here, $I=12\neq 1112=J$, and we have highlighted in blue the vertices corresponding to the bounded dimension $r_{N}$. Recall that edges without arrows correspond to the matrix $\mathbf{I}$ (matrix of ones), and that two edges corresponding to matrices $A$ and $B$ which share direction and terminal vertices can be merged into the edge $A\odot B$.
Appendix E Experiment setup
State tracking.
We train all models for $5$ epochs, with a batch size of $128$, $3$ different random seeds, learning rate set to $0.0001$, weight decay set to $0.01$, gradient clipping 1.0, and the AdamW optimizer (Loshchilov & Hutter, 2017). We sample $1.6$M samples from all the possible permutations, and split the data with a ratio of 4 to 1. We use the implementation and the hyperparameters provided by Merrill et al. (2024). We train the model for sequence length $16$, and evaluate for sequence length $2$ through $50$.
Copying.
Mod arithmetic.

### v2

@@ Appendix F Low-Rank Expressiveness
Following the notation and methodology of Cirone et al. (2024b)[B.4] ), this can be written in terms of the Signature as
@@
It follows directly from Eq. 19 that any linear readout of $Y_{t}$ can be represented as a series in signature terms. As a result, these systems are fundamentally restricted to learning functions that closely approximate these convergent series.
@@
In our setting $\frac{1}{N}\langle A_{I}Y_{0},A_{J}Y_{0}\rangle_{\mathbb{R}^{N}}$ corresponds to a product graph $G_{I,J}$ corresponding to a ladder having $2|I|+2|J|$ edges as shown in figure 12. We can then use (Cirone et al., 2024a)[Prop. 2] to compute the square of the $L^{2}$ norm in equation Eq. 21, the only difference from the dense case is that half of the vertices (excluding the "middle" one) correspond to a space of dimension $r_{N}$ while the rest to the standard $N$.
@@
Since the total number of admissible pairings of $G_{I,J}\sqcup G_{I,J}$ is $\bigl{(}4(|I|+|J|)\bigr{)}!!$!, we conclude that equation 21 holds with $\kappa=4$ and $o(1):=\mathcal{O}\bigl{(}\frac{1}{\sqrt{N\wedge r_{N}}}\bigr{)}$.
@@
Remark F.4.
While the proof crucially uses the assumption $r_{N}\to\infty$ as $N\to\infty$, at the same time we have not provided an argument against $r_{N}$ not diverging. In figure 13 we present a counterexample, showing that if $r_{N}$ does not diverge then the asymptotics differ from the dense ones, in particular some symmetries are "lost", impossible to recover due to unavoidable noise.
Figure 13: Admissible pairing different from the "identity" paring, but still leading to maximal asymptotic scaling in the bounded $r_{N}$ case. Here, $I=12\neq 1112=J$, and we have highlighted in blue the vertices corresponding to the bounded dimension $r_{N}$. Recall that edges without arrows correspond to the matrix $\mathbf{I}$ (matrix of ones), and that two edges corresponding to matrices $A$ and $B$ which share direction and terminal vertices can be merged into the edge $A\odot B$.